rearrange
5(2x-y)=3y-6
to find y in terms of x

Answer by ewatrrr(24785)

5(2x-y)=3y-6
10x - 5y = 3y - 6
10x + 6 = 8y
(10/8)x + 6/8 = y
y = (5/4)x + 3/4